### v1

@@ 2.3 Markov kernels and stability of the optimal filter
Let $\{\Phi_{n}\}_{n\geq 0}$ be the sequence of PU operators generated by the Markov kernels $K=\{K_{n}\}_{n\geq 1}$ and the likelihoods $g=\{g_{n}\}_{n\geq 1}$ according to (3).
@@
The terminology ‘stability of the optimal filter’ is well established in the literature [DelMoral01c, DelMoral04]. We note, however, that stability is a property of the operators $\Phi_{n}$, rather than the probability laws $\pi_{n}$ (see also [Crisan20] for further details).
A sufficient condition for $\Phi_{n}$ to be stable is that the Markov kernels in $K$ are mixing and the likelihood functions in $g$ integrable w.r.t. some reference probability density functions (pdfs) $u=\{u_{n}\}_{n\geq 1}$ on $\mathcal{X}$}.
@@
Then, for any $n\geq m\geq 0$, the composition of PU operators satisfies the inequality
@@
2.4 Particle filters
A particle filter is a recursive Monte Carlo method to approximate the sequence of measures $\{\pi_{n}\}_{n\geq 0}$. Its simplest version, often termed the bootstrap filter or the standard particle filter [Gordon93, Kitagawa96] (see also [Liu98, Doucet00, Djuric03]), is outlined in Algorithm 1. The Monte Carlo estimate of $\pi_{n}$ is denoted $\pi_{n}^{N}$ and can be written as
(7) $\pi_{n}^{N}({\sf d}x)=\frac{1}{N}\sum_{i=1}^{N}\delta_{x_{n}^{i}}({\sf d}x),$
where $\delta_{x}$ is the Dirac delta measure located at $x\in{\mathcal{X}}_{n}$ and $x_{n}^{1},\ldots,x_{n}^{N}$ are Monte Carlo samples from the state space $\mathcal{X}$}.
Algorithm 1 The standard particle filter for a SSM ${\mathcal{S}}=(\pi_{0},K,g)$.
  • Inputs: model ${\mathcal{S}}=(\pi_{0},K,g)$ and number of particles $N$.
    Outputs: approximation $\pi_{n}^{N}$ of the optimal filter $\pi_{n}$ at every step $n\geq 1$.
  • Initialisation. At time $n=0$, draw $N$ iid samples $x_{0}^{i}\sim\pi_{0}$, $i=1,\ldots,N$.
  • Recursive step. At time $n\geq 1$:
    1. 1.
      For $i=1,\ldots,N$,
      • draw $\bar{x}_{n}^{i}\sim K_{n}(x_{n-1},{\sf d}x)$ and
      • compute $\bar{w}_{n}^{i}=g_{n}\left(\bar{x}_{n}^{i}\right)$.
    2. 2.
      Normalise the weights, $w_{n}^{i}=\frac{\bar{w}_{n}^{i}}{\sum_{j=1}^{N}\bar{w}_{n}^{j}}$, $i=1,\ldots,N$.
    3. 3.
      Resample: let $\bar{\pi}_{n}^{N}({\sf d}x)=\sum_{j=1}^{N}w_{n}^{j}\delta_{\bar{x}_{n}^{j}}({\sf d}x)$ and draw $N$ iid samples $x_{n}^{i}\sim\bar{\pi}_{n}^{N}$, $i=1,\ldots,N$.
If the Markov SSM of interest, ${\mathcal{S}}=(\pi_{0},K,g)$, yields a stable optimal filter, then it can be proved that the particle filter approximation $\pi_{n}^{N}$ converges uniformly over time towards the optimal filter $\pi_{n}$. In particular, one has the following result.
Theorem 2.6.
Let ${\mathcal{S}}=(\pi_{0},K,g)$ be a Markov SSM and assume that
  • (a)
    the kernels $K=\{K_{n}\}_{n\geq 1}$ are mixing with constant $\gamma>0$, and
  • (b)
    there are uniform bounds $\inf_{n\geq 1,x\in{\mathcal{X}}}g_{n}(x)>\zeta>0$ and $\|g\|_{\infty}:=\sup_{n\geq 1}\|g_{n}\|_{\infty}<\infty$ on the potentials $g=\{g_{n}\}_{n\geq 1}$
Then, for any measurable $f:{\mathcal{X}}\mapsto\mathbb{R}$ with $\|f\|\leq 1$, there is a constant $c_{p}<\infty$, independent of $n$, NN, $f$, $d_{x}$ and $d_{y}$, such that
$\sup_{n\geq 1}\|\pi_{n}(f)-\pi_{n}^{N}(f)\|_{p}\leq\frac{c_{p}\|g\|_{\infty}}{\zeta\gamma^{6}}\times\frac{1}{\sqrt{N}}\quad\text{for all $p\geq 1$.}.$
Uniform convergence of the PF is also well known and similar results can be found in [DelMoral04, Kunsch05, Whiteley13b]. We simply note that the mixing assumption (a) in Theorem 2.6 is the same as in [Kunsch05, Theorem 2] (which states uniform convergence in probability of $\pi_{n}^{N}\to\pi_{n}$), and slightly weaker than the usual mixing assumption in [DelMoral04, Theorem 7.4.4]. An explicit proof is presented in Appendix A.
@@ 3.1 Constrained models
(8) $\hat{K}_{n}(x^{\prime},{\sf d}x):=\frac{\mathbbm{1}_{{\mathcal{C}}_{n}}(x)K_{n}(x^{\prime},{\sf d}x)}{\int_{{\mathcal{C}}_{n}}K_{n}(x^{\prime},{\sf d}x)},$
$\mathbbm{1}_{{\mathcal{C}}_{n}}$ is the indicator function for the set in the subscript, and $\hat{\pi}_{0}({\sf d}x)=\frac{\mathbbm{1}_{{\mathcal{C}}_{0}}(x)\pi_{0}({\sf d}x)}{\int_{{\mathcal{C}}_{0}}\pi_{0}({\sf d}x)}$. It is easy to see that $\hat{K}_{n}(x^{\prime},{\mathcal{X}})=1$ for any $x^{\prime}\in{\mathcal{X}}$ and hence the kernels in $\hat{K}$ are well defined whenever the original kernels in $K$ are well defined.
The PU operator for model $\hat{\mathcal{S}}$ is constructed as in Eq. (3), i.e., for a real, measurable test function $f:{\mathcal{X}}\mapsto\mathbb{R}$ we have
@@ 3.2 Stability
Let us assume that the SSM of interest, $\mathcal{S}$}, satisfies some mild regularity conditions regarding the Markov kernels $K=\{K_{n}\}_{n\geq 1}$ and the potentials $g=\{g_{n}\}_{n\geq 1}$.
@@
where the densities $k_{n}(x^{\prime},x)$ and $k_{n}(x_{o},x)$ are not normalised on ${\mathcal{C}}_{n}$. Moreover, (10) also yields
@@
where $\hat{k}_{n}(x^{\prime},x)=\frac{k_{n}(x^{\prime},x)}{\int_{{\mathcal{C}}_{n}}K_{n}(x^{\prime},{\sf d}x)}$ and $u_{n}(x)=\frac{k_{n}(x_{o},x)}{\int_{{\mathcal{C}}_{n}}K_{n}(x_{o},{\sf d}x)}$ are normalised probability densities on ${\mathcal{C}}_{n}$. The inequalities in (14) imply that the constrained kernels $\hat{K}=\{\hat{K}_{n}\}_{n\geq 1}$ are mixing with constant $\gamma:=\frac{\bar{\gamma}^{2}}{\|k\|_{\infty}^{2}}>0$ and the pdfs $u_{n}(x)$ satisfy $\int_{{\mathcal{C}}_{n}}g_{n}(x)u_{n}(x){\sf d}x<\|g\|_{\infty}<\infty$.
@@
If the state $X_{n}$ is a homogeneous process, i.e., $K_{n}=K$ for some Markov kernel $K:{\mathcal{X}}\times{\mathcal{B}}({\mathcal{X}})\mapsto[0,1]$, and Assumption 3.2 holds, it is always possible to choose compact sets $\{{\mathcal{C}}_{n}\}_{n\geq 1}$ such that the uniform lower bound in (10) exists. Also note that, while Gaussian kernels are not mixing according to Definition 2.4, they satisfy the conditions in Assumption 3.2 and, therefore, they can be constrained to yield a stable sequence of posterior measures $\hat{\pi}_{n}$, $n\geq 0$.
@@ 3.3 Approximation errors
We are interested in constructing constrained models $\hat{\mathcal{S}}$ for which the posterior marginals $\hat{\pi}_{n}$ are close to the marginals $\pi_{n}$ generated by the original model $\mathcal{S}$} under (reasonable) regularity assumptions. To this end, we adopt the framework of [Crisan20]. In particular, consider the sequence $\{{\mathcal{C}}^{l}\}_{l\geq 1}$ where each element ${\mathcal{C}}^{l}=\{{\mathcal{C}}_{n}^{l}\}_{n\geq 0}$ is a family of compact sets that can be used to constrain the model $\mathcal{S}$}. Specifically, we construct the sequence of constrained models $\hat{\mathcal{S}}^{l}=(\hat{\pi}_{0}^{l},\hat{K}^{l},g)$, where $\hat{K}^{l}=\{\hat{K}_{n}^{l}\}_{n\geq 1}$,
@@
The family of compact subsets $\{{\mathcal{C}}_{n}^{l}\}_{n\geq 0,l\geq 1}$ of the state space $\mathcal{X}$} is constructed to guarantee that, for every $n\geq 0$, $\lim_{l\to\infty}{\mathcal{C}}_{n}^{l}={\mathcal{X}}$ and there are positive sequences $\{\varepsilon_{n}^{l}\}_{n\geq 0,l\geq 1}$, increasing with $l$, such that
@@
See Appendix B for a proof. The terms $\varepsilon_{i}^{l}$, $i\leq n$, are the increasing sequences in Assumption 3.3, which satisfy $\lim_{l\to\infty}\varepsilon_{i}^{l}=1$, hence $\left|\pi_{n}(f)-\hat{\pi}_{n}^{l}(f)\right|\to 0$ as expected from Lemma 2.2.
@@
$\pi_{0}({\mathcal{C}}_{0}^{l})\geq\varepsilon^{l},\quad\inf_{x^{\prime}\in{\mathcal{C}}_{n-1}^{l}}\int_{{\mathcal{C}}_{n}^{l}}K_{n}(x^{\prime},{\sf d}x)\geq\varepsilon^{l}\quad\text{and}\quad\lim_{l\to\infty}\varepsilon^{l}=1~~\forall n\geq 0.$
This stronger assumption on the constraints ${\mathcal{C}}=\{{\mathcal{C}}^{l}\}_{l\geq 1}$ yields a straightforward corollary of Theorem 3.3.
@@
for any pair of probability measures $\mu,\eta\in{\mathcal{P}}({\mathcal{X}})$.
Under Assumption 3.3 it is possible to obtain uniform (over time $n$) error bounds on the approximation error $|\pi_{n}(f)-\hat{\pi}_{n}^{l}(f)|$. However, similar to Theorem 2.6, the attainment of uniform bounds demands some additional conditions related to the likelihoods $g_{n}$. Specifically, we have worked under two alternative assumptions:
  • the likelihoods are uniformly bounded away from 0, i.e., $\inf_{n,x}g_{n}(x)>\zeta>0$, as in the assumptions of Theorem 2.6 (or in Theorem 7.4.4 in [DelMoral04]), or
  • the likelihoods are continuous, vanishing as $\lim_{\|x\|\to\infty}g_{n}(x)=0$, and the compact sets ${\mathcal{C}}_{n}^{l}$ are chosen to be superlevel sets of $g_{n}$ for every $n\geq 1$, i.e., ${\mathcal{C}}_{n}^{l}=\{x\in{\mathcal{X}}:g_{n}(x)\geq\zeta_{n}^{l}\}$ for some real sequences $\zeta_{n}^{l}>0$, with $\lim_{l\to\infty}\zeta_{n}^{l}=0$ for all $n$.
With the first assumption, we have Theorem 3.5 below, while the second assumption yields Theorem 3.6.
@@
See Appendix C for a proof. Note that the supremum in (15) is taken over all real and measurable test functions $f:{\mathcal{X}}\mapsto\mathbb{R}$ such that $\|f\|_{\infty}\leq 1$.
@@
We also note that, for many models, Assumption 3.3 (which is used in both Theorems 15 and 16) can be hard to reconcile with the the requirement that the compact sets ${\mathcal{C}}_{n}^{l}$ in the constraint be all of them level supersets of their respective likelihoods $g_{n}$. However, the latter condition is appealing because it yields a practical guideline for the construction of constraints and avoids the need for a uniform positive lower bound on the likelihoods (a condition that does not hold for many models). If we impose the weaker Assumption 3.3 on the family of constraints $\{{\mathcal{C}}^{l}\}_{l\geq 1}$, then the error bounds are not uniform over time any more, however we can still have the following result.
@@
3.4 Constrained particle filters
In this section we investigate to what extent the standard particle filter approximation $\hat{\pi}_{n}^{l,N}$ of the constrained marginal law $\hat{\pi}_{n}^{l}$ can be used as an estimate of the optimal (unconstrained) Bayesian filter $\pi_{n}$. This can be done in different settings, however we adopt the key assumption $\inf_{n\geq 1}g_{n}>\zeta>0$ in order to provide error bounds comparable to those in Theorem 2.6.
Theorem 3.9.
Let ${\mathcal{S}}=(\pi_{0},K,g)$ be a SSM for which Assumptions 3.2 and 3.3 hold and, in addition, there is a constant $\zeta>0$ such that $\inf_{n\geq 1}g_{n}>\zeta$. Let $\{{\mathcal{C}}^{l}\}_{l\geq 1}$ be a sequence of constraints of the state space $\mathcal{X}$} that satisfies Assumption 3.3 and the inequality
(17) $\inf_{n\geq 1,(x^{\prime},x)\in{\mathcal{C}}_{n-1}^{l}\times{\mathcal{C}}_{n}^{l}}k_{n}(x^{\prime},x)\geq\bar{\gamma}_{l}>0.$
Then, for any $p\geq 1$,
(18) $\sup_{n\geq 1,|f|\leq 1}\|\pi_{n}(f)-\hat{\pi}_{n}^{l,N}(f)\|_{p}<c\frac{1-\varepsilon^{l}}{\varepsilon^{l}}+\frac{c_{p}\|g\|_{\infty}}{\zeta\gamma_{l}^{6}\sqrt{N}},$
where $c_{p}<\infty$ is a constant that depends only on $p$. In particular, for every $l<\infty$,
(19) $\lim_{N\to\infty}\sup_{n\geq 1,|f|\leq 1}\|\pi_{n}(f)-\hat{\pi}_{n}^{l,N}(f)\|_{p}<c\frac{1-\varepsilon^{l}}{\varepsilon^{l}}$
and there is some constant $\bar{c}<\infty$, independent of $N$ and $n$, such that
(20) $\sup_{n\geq 0,|f|\leq 1}\lim_{l\to\infty}\|\pi_{n}(f)-\hat{\pi}_{n}^{l,N}(f)\|_{p}<\frac{\bar{c}}{\sqrt{N}}.$
A triangle inequality yields $\|\pi_{n}(f)-\hat{\pi}_{n}^{l,N}(f)\|_{p}\leq\|\pi_{n}(f)-\hat{\pi}_{n}^{l}(f)\|_{p}+\|\hat{\pi}_{n}^{l}(f)-\hat{\pi}_{n}^{l,N}(f)\|_{p}$ and we bound the two terms on the right-hand side separately. In particular, the first term on the right-hand-side of inequality (18) follows from Theorem 3.5, while the second term is obtained via Theorem 3.1 (which ensures the exponential stability of the constrained operators $\hat{\Phi}_{n}^{l}$) together with Theorem 2.6 (for the uniform convergence over time of $\hat{\pi}_{n}^{l,N}$ towards $\hat{\pi}_{n}^{l}$). The limit (19) is straightforward, while he bound in (20) can be obtained by the same argument as in the proof of Theorem 2.6. See Appendix F for a complete proof.
3.5 High-dimensional models
The dependence of the potential functions $g_{n}(x)$ on the dimension $d_{y}$ of the observation space $\mathcal{Y}$} has been found to be a major cause for the “collapse of the particle filter” [Bengtsson08, Snyder08] when $d_{x}\geq d_{y}>>1$, as it causes the importance weights to degenerate, meaning that $\max_{1\leq i\leq n}w_{n}^{i}\to 1$ as $d_{y}\to\infty$. This phenomenon can be alleviated within the framework described in Sections 3.3 and 3.4.
To be specific, choose some functions $\zeta_{n}:\mathbb{N}\mapsto(0,\|g_{n}\|_{\infty}]$ and take superlevel sets ${\mathcal{C}}=\{{\mathcal{C}}_{n}\}_{n\geq 0}$ such that, for every $n\geq 1$, ${\mathcal{C}}_{n}=\{x\in{\mathcal{X}}:g_{n}(x)\geq\zeta_{n}(d_{y})\}$. Then consider the (normalised) effective sample size (ESS) estimator [Kong94, Doucet00]
$\text{ESS}_{N}:=\frac{1}{N\sum_{i=1}^{N}\left(w_{n}^{i}\right)^{2}}$
which is usually employed to assess whether the weights are evenly distributed or, on the contrary, one (or a few) particles accumulate all the weight. It can be easily seen that $\frac{1}{N}\leq\text{ESS}_{N}\leq 1$: if the particle weights are uniform, $w_{n}^{i}=\frac{1}{N}$ (which would be the case when sampling directly from the optimal filter), then $\text{ESS}_{N}=1$, while $\text{ESS}_{N}=\frac{1}{N}$ when $\max_{1\leq i\leq N}w_{n}^{i}=1$, i.e., there is a single particle that accumulates all the weight. A low (close to $\frac{1}{N}$) $\text{ESS}_{N}$ indicates that the particle filter has degenerated [Kong94, Doucet00] (or collapsed in the terminology of [Snyder08, Bengtsson08]). It can be seen that, for the constrained particle filter,
$\text{ESS}_{N}\geq\frac{\sup_{x\in{\mathcal{C}}_{n}}g_{n}(x)}{\inf_{x\in{\mathcal{C}}_{n}}g_{n}(x)}\geq\frac{\|g_{n}\|_{\infty}}{\zeta_{n}(d_{y})},$
hence the construction of the sets ${\mathcal{C}}_{n}$ in the constraint ${\mathcal{C}}=\{{\mathcal{C}}_{n}\}_{n\geq 0}$ as superlevel sets of $g_{n}$ provides a way to control degeneracy of the weights (possibly at the cost of an increased approximation error, depending on the model).

### v2

@@ 2.3 Markov kernels and stability of the optimal filter
Let $\{\Phi_{n}\}_{n\geq 0}$ be the sequence of PU operators generated by the Markov kernels $K=\{K_{n}\}_{n\geq 1}$ and the likelihoods $g=\{g_{n}\}_{n\geq 1}$ according to (3) (and note that $\Phi_{n}$ does not depend on $\pi_{0}$).
@@
The terminology “stability of the optimal filter” is well established in the literature [DelMoral01c, DelMoral04]. We note, however, that stability is a property of the operators $\Phi_{n}$, rather than the probability laws $\pi_{n}$ (see also [Crisan20] for further details). A sufficient condition for $\Phi_{n}$ to be stable is that the Markov kernels in $K$ are mixing and the likelihood functions in $g$ integrable w.r.t. some reference probability density functions (pdfs) $u=\{u_{n}\}_{n\geq 1}$ on $\mathcal{X}$}.
@@
Then, for any $n>m\geq 0$, the composition of PU operators satisfies the inequality
@@ 3.1 Constrained models
(7) $\hat{K}_{n}(x^{\prime},{\sf d}x):=\frac{\mathbbm{1}_{{\mathcal{C}}_{n}}(x)K_{n}(x^{\prime},{\sf d}x)}{\int_{{\mathcal{C}}_{n}}K_{n}(x^{\prime},{\sf d}\tilde{x})}$,
$\mathbbm{1}_{{\mathcal{C}}_{n}}$ is the indicator function for the set in the subscript, and $\hat{\pi}_{0}({\sf d}x)=\frac{\mathbbm{1}_{{\mathcal{C}}_{0}}(x)\pi_{0}({\sf d}x)}{\int_{{\mathcal{C}}_{0}}\pi_{0}({\sf d}x)}$. It is easy to see that $\int_{{\mathcal{C}}_{n}}\hat{K}_{n}(x^{\prime},{\sf d}x)=1$ for any $x^{\prime}\in{\mathcal{X}}$ and hence the kernels in $\hat{K}$ are well defined.
The PU operator $\hat{\Phi}_{n}$ for model $\hat{\mathcal{S}}$ is constructed as in Eq. (3), i.e., for a real, measurable test function $f:{\mathcal{X}}\mapsto\mathbb{R}$ we have
@@ 3.2 Stability
Let us assume that the SSM of interest, $\mathcal{S}$}, satisfies some mild regularity conditions regarding the Markov kernels $K=\{K_{n}\}_{n\geq 1}$ and the likelihoods $g=\{g_{n}\}_{n\geq 1}$.
@@
where the densities $k_{n}(x^{\prime},x)$ and $k_{n}(x_{o},x)$ are not normalized on ${\mathcal{C}}_{n}$. Moreover, (9) and Assumption 3.2 also yield
@@
where $\hat{k}_{n}(x^{\prime},x)=\frac{k_{n}(x^{\prime},x)}{\int_{{\mathcal{C}}_{n}}K_{n}(x^{\prime},{\sf d}x)}$ and $u_{n}(x)=\frac{k_{n}(x_{o},x)}{\int_{{\mathcal{C}}_{n}}K_{n}(x_{o},{\sf d}x)}$ are normalized probability densities on ${\mathcal{C}}_{n}$. The inequalities in (13) imply that the constrained kernels $\hat{K}=\{\hat{K}_{n}\}_{n\geq 1}$ are mixing with constant $\gamma:=\frac{\bar{\gamma}^{2}}{\|k\|_{\infty}^{2}}>0$ and the pdfs $u_{n}(x)$ satisfy $\int_{{\mathcal{C}}_{n}}g_{n}(x)u_{n}(x){\sf d}x<\|g\|_{\infty}<\infty$.
@@
If the state $X_{n}$ is a homogeneous process, i.e., $K_{n}=K$ for some Markov kernel $K:{\mathcal{X}}\times{\mathcal{B}}({\mathcal{X}})\mapsto[0,1]$, and Assumption 3.2 holds, it is always possible to choose compact sets $\{{\mathcal{C}}_{n}\}_{n\geq 1}$ such that the uniform lower bound in (9) exists. Also note that, while Gaussian kernels are not mixing according to Definition 2.4, they satisfy Assumption 3.2 and, therefore, they can be constrained to yield a stable sequence of posterior measures $\hat{\pi}_{n}$, $n\geq 0$.
@@ 3.3 Approximation errors
We are interested in constructing constrained models $\hat{\mathcal{S}}$ for which the posterior marginals $\hat{\pi}_{n}$ are close to the marginals $\pi_{n}$ generated by the original model $\mathcal{S}$} under (reasonable) regularity assumptions. To this end, we adopt the framework of [Crisan20]. In particular, consider the sequence $\{{\mathcal{C}}^{l}\}_{l\geq 1}$ where each element ${\mathcal{C}}^{l}=\{{\mathcal{C}}_{n}^{l}\}_{n\geq 0}$ is a sequence of compact sets that can be used to constrain the model $\mathcal{S}$}. Specifically, we construct the sequence of constrained models $\hat{\mathcal{S}}^{l}=(\hat{\pi}_{0}^{l},\hat{K}^{l},g)$, where $\hat{K}^{l}=\{\hat{K}_{n}^{l}\}_{n\geq 1}$,
@@
The family of compact subsets $\{{\mathcal{C}}_{n}^{l}\}_{n\geq 0,l\geq 1}$ of the state space $\mathcal{X}$} is constructed to guarantee that, for every $n\geq 0$, $\lim_{l\to\infty}{\mathcal{C}}_{n}^{l}={\mathcal{X}}$ and there are positive sequences $\{\varepsilon_{n}^{l}\}_{n\geq 0,l\geq 1}$, increasing with $l$ for every fixed $n$, such that
@@
See Appendix A for a proof. The terms $\varepsilon_{i}^{l}$, $i\leq n$, are the increasing sequences in Assumption 3.3, which satisfy $\lim_{l\to\infty}\varepsilon_{i}^{l}=1$, hence $\left|\pi_{n}(f)-\hat{\pi}_{n}^{l}(f)\right|\to 0$ as expected from Lemma 2.2.
@@
$\pi_{0}({\mathcal{C}}_{0}^{l})\geq\varepsilon^{l},\quad\inf_{x^{\prime}\in{\mathcal{C}}_{n-1}^{l}}\int_{{\mathcal{C}}_{n}^{l}}K_{n}(x^{\prime},{\sf d}x)\geq\varepsilon^{l}\quad\text{and}\quad\lim_{l\to\infty}\varepsilon^{l}=1.$
This stronger assumption on the constraints $\{{\mathcal{C}}^{l}\}_{l\geq 1}$ yields a straightforward corollary of Theorem 3.3.
@@
for any pair of probability measures $\mu,\eta\in{\mathcal{P}}({\mathcal{X}})$ and $n\geq m$.
Under Assumption 3.3 it is possible to obtain uniform (over time $n$) bounds on the approximation error $|\pi_{n}(f)-\hat{\pi}_{n}^{l}(f)|$. However, the attainment of uniform approximation rates demands some additional conditions related to the likelihoods $g_{n}$. Specifically, we have worked under two alternative assumptions:
  • (a)
    the likelihoods are uniformly bounded away from 0, i.e., $\inf_{n,x}g_{n}(x)>\zeta>0$, or
  • (b)
    the likelihoods are continuous, vanishing as $\lim_{\|x\|\to\infty}g_{n}(x)=0$, and the compact sets ${\mathcal{C}}_{n}^{l}$ are superlevel sets of $g_{n}$ for every $n\geq 1$, i.e., ${\mathcal{C}}_{n}^{l}=\{x\in{\mathcal{X}}:g_{n}(x)\geq\zeta_{n}^{l}\}$ for some real sequences $\zeta_{n}^{l}>0$, with $\lim_{l\to\infty}\zeta_{n}^{l}=0$ for all $n$.
Assumptions (a) and (b) above yield Theorem 3.5 and Theorem 3.6, respectively.
@@
See Appendix B for a proof. Note that the supremum in (14) is taken over all real and measurable test functions $f:{\mathcal{X}}\mapsto\mathbb{R}$ such that $\|f\|_{\infty}\leq 1$.
@@
We also note that, for many models, Assumption 3.3 (which is used in both Theorems 14 and 15) can be hard to reconcile with the the requirement that the compact sets ${\mathcal{C}}_{n}^{l}$ in the constraint be all of them superlevel sets of their respective likelihoods $g_{n}$. However, the latter condition is appealing because it yields a practical guideline for the construction of constraints and avoids the need for a uniform positive lower bound on the likelihoods (a condition that does not hold for many models). If we impose the weaker Assumption 3.3 on the family of constraints $\{{\mathcal{C}}^{l}\}_{l\geq 1}$, then the error bounds are not uniform over time any more, however we can still have the following result.